# Towards Wireless-Native Big AI Model: Insights into Its Ambitions, Peculiarities and Methodologies

### v2

@@ Abstract
Researches on leveraging big artificial intelligence model (BAIM) technology to drive the intelligent evolution of wireless networks are emerging. However, since the breakthrough in generalization brought about by BAIM techniques mainly occurs in natural language processing, there is still a lack of a clear technical roadmap on how to efficiently apply BAIM techniques to wireless systems with many additional peculiarities. To this end, this paper first reviews recent research works on BAIM for wireless and assesses the current research situation. Then, this paper analyzes and compares the differences between language intelligence and wireless intelligence on multiple levels, including scientific foundations, core usages, and technical details. It highlights the necessity and scientific significance of developing BAIM technology in a wireless-native way, as well as new issues that need to be considered in specific technical implementation. Finally, by synthesizing the evolutionary laws of language models with the particularities of wireless system, this paper provides several instructive methodologies for how to develop wireless-native BAIM.
@@ 1 Introduction
Given the outstanding potential of AI technology in complex feature extraction, high-dimensional data representation and adaptive decision-making, how to apply AI technology to solve wireless problems has been a critical topic in wireless technology research [1, 2, 3, 4, 5, 6]. However, simple migrating or reusing successful approaches in computer vision (CV) or natural language processing (NLP) to wireless systems tends to face serious bottlenecks in terms of performance and generalizability [7]. This “incompatibility” is due to the fact that AI for wireless is not directly equivalent to AI for natural language or computer vision. Specifically, AI for wireless essentially requires to learn electromagnetic physics and engineering methods, which makes it quite different from modeling human cognition and interaction capabilities in NLP and CV. Therefore, high-quality wireless AI technologies need to adapt to the characteristics of wireless data and the unique requirements of wireless systems, a concept known as ‘wireless-native’. Recent years, the incorporation of wireless-native methods such as physics-inspired structure [7, 8, 9, 10, 11, 12], and cross-module combination [13, 14] has significantly improved the performance and generalization of wireless AI. Meanwhile, the introduction of prior information and advanced learning architectures can also reduce the demand for training data and computational resources. These fundamental advantages are of great significance for advancing the usability of wireless AI, whether in centralized or distributed usage scenarios [2].
@@
Although much progress has been made in wireless AI research, it still only produces intelligence limited to specific tasks and specific scenarios, which has obvious weaknesses in generalizability and stability. While recently in NLP field, large language models (LLMs) have demonstrated the outstanding potential of big AI models (BAIMs) in board generalization and high reliability. With the help of a pre-trained LLM, numerous language tasks such as translation, question&answer (Q&A) and summarization can be accomplished with state-of-the-art performance. This excellent generalization is exactly what wireless AI desperately needs today. Timely, researchers in [15] first defined wireless-native BAIM (wBAIM). wBAIM takes integrating multi-tasks, unifying multi-scenarios and all-in-one scheduling as its typical features, and emphasizes developing native BAIM focusing on the physical nature of wireless data and the fundamental requirements of wireless systems. Realizing these promising visions represents a quantum leap in functionality, universality, and reliability for wireless AI, but it will also require comprehensive and in-depth technical exploration as support. Although [15] has summarized the technical challenges and potential solutions of wBAIM, from the recent research trends in integrating BAIM and wireless, there are several issues about wBAIM that urgently need to be clarified and supplemented as follows (the graphic illustration is shown in Fig. 1):
@@ 1. $\bullet$
    Question 1: What are the necessity and significance of investigating native BAIM for wireless in the context of powerful LLMs and their excellent transferability?
@@ 2. $\bullet$
    Question 2: What are the technical peculiarities of wBAIM, especially in comparison to conventional wireless models and existing LLMs?
@@ 3. $\bullet$
    Question 3: What methodologies for developing wBAIM can be derived by synthesizing the evolutionary laws of language models and the peculiarities revealed in Question 2?
@@
To answer these questions, this paper is organized as follows. Section 2 provides a review of relevant advancements, providing an overview of the current research landscape. Next, in Section 3, the paper analyzes the fundamental differences between language and wireless, highlighting the necessity, challenges, and significant value of studying wBAIM. Section 4 then examines, from a system-level perspective, the unique technical considerations essential for wBAIM research. After that, Section 5 presents several forward-looking insights into how to develop wBAIM. Finally, the conclusions are drawn in Section 6.
@@ 2 Literature Review
The success of LLMs have motivated numerous researches on BAIM for wireless to accelerate wireless systems towards ubiquitous intelligence. Existing researches can be generally categorized into the following three research paradigms: wireless LLM, LLM-based wireless agent, and wBAIM. The overall architectures of these three paradigms are shown in Fig. 2.
@@ 2.1 Wireless LLM
Wireless LLM refers to applying existing LLMs to wireless problems through reuse, fine-tuning, or retrieval augmented generation (RAG) techniques, etc. These methods translate requirements in wireless systems into language-based description, positioning them within or approximately within the generalization capabilities of LLMs. Leveraging the advanced information comprehension and vast knowledge base of LLMs, initial successes have been achieved in protocol understanding, command interaction, and resource allocation in wireless systems [16, 17]. Moreover, some researchers have applied LLMs to NLP-related tasks in wireless systems, such as text categorization [18], text summarization, and question answering [19, 20], with specific applications including AI platforms for customer interaction, virtual assistants for technicians, network diagnostic management, and telecom standards question&answer bots [21, 22, 23]. While these studies demonstrate the value of LLMs in wireless systems, they primarily remain within the scope of traditional NLP. Furthermore, LLMs have been applied to practical requirements in wireless systems such as power control, intrusion detection, resource allocation, and symbol demodulation [24, 25, 26, 27]. By constructing appropriate prompts, the LLM based on in-context learning can understand the wireless problems to a certain extent, thus enabling initial wireless-specific functions without the need for parameter adjustments.
Meanwhile, to further enhance the applicability of LLMs in wireless communications, some researchers have begun exploring the integration of wireless domain knowledge into LLMs and focusing on the role of multimodal information [28]. These works first pre-train generative models on specific telecom knowledge datasets, followed by supervised fine-tuning (SFT) of the models with telecom instruction sets [29] and aligning the model’s outputs with human preferences through reinforcement learning from human feedback (RLHF) [30, 31, 32]. In addition to direct re-pretraining, the RAG mechanism and the Langchain framework have also been identified as important approaches to supplement the telecom knowledge base [33, 34, 35, 36, 37]. Additionally, some researchers have constructed foundational telecom text datasets to serve as materials for expanding wireless knowledge in LLMs. These include SPEC5G, a text categorization dataset derived from 3GPP technical reports [38]; TeleQnA, a Q&A dataset extracted from technical documents and research materials [39]; TSpec-LLM, a dataset based on 3GPP specifications that preserves table formulas [40]; Tele-Data, a LaTeX-formatted dataset obtained from sources like arXiv, 3GPP standards, and Wikipedia; and Tele-Eval, a Q&A evaluation dataset [41].
@@ 2.2 LLM-Based Wireless Agent
Although LLM-based methods have been somewhat explored, the current technical progress primarily focuses on instruction design, resource scheduling, and a few simple wireless problems. These approaches still lack effective solutions for numerous tough wireless challenges, such as scenario sensing and high-rate transmission. This limitation arises from two primary factors: firstly, linguistic expressions are frequently insufficient in precisely characterizing multimodal signals and their intricate correlations; secondly, the language intelligence derived from textual training lacks adequate knowledge reserves to comprehend the data structures and underlying mechanisms inherent in wireless data. Therefore, instead of embracing LLMs, some researchers have explored BAIM in a wireless-native way. Wireless-native AI refers to an intelligent framework based on regular patterns of AI technologies, such as deep neural networks, while incorporating the peculiarities of wireless systems. It tailors model design to the nature of wireless data and system requirements, and adapts through re-learning on wireless data. Conventional wireless AI has also followed this approach in many studies. For example, in model design, researchers initially leveraged the numerical characteristics of wireless data, such as exploiting channel sparsity in the angle-delay domain and transforming it into the angle domain to make it more image-like [47, 48]. Later, researchers delved deeper into the physical mechanisms underlying the numerical characteristics and employed more fundamental, physics-inspired techniques to guide model optimization [7, 8, 9, 10, 12]. In terms of architecture, some cross-module [13, 14] and cross-layer [49] joint techniques have also been developed to facilitate applications in wireless systems.
However, given that wBAIM has a broader functional scope, relying solely on existing wireless-native AI technologies is insufficient. Therefore, in [15], the authors first outlined three key generalizability indicators of wBAIM: multi-task integration, multi-scenario unification, and all-in-one scheduling. Building on this, [50] introduced SWTCAN, a network based on the advanced swin transformer, which integrates downlink pilot patterns, channel estimation, and uplink channel feedback into a joint design. [51] proposed Csi-LLM, a downlink channel prediction model based on GPT2, which supports variable-length inputs and enables continuous prediction. Similarly, in [52], a pre-training and fine-tuning based channel prediction method was presented, enhancing performance in few-shot learning. Meanwhile, [53] evaluated the similarity and diversity across multi-scenario datasets, offering valuable insights into developing multi-scenario generalization techniques from a data-based perspective.
@@ 2.4 Summary of the literature
Fig. 3 presents statistical information on the 36 publications related to BAIM in wireless systems included in this paper. It shows that research interest in BAIM for wireless is growing, and there is a growing consensus on the importance of BAIM techniques for the evolution of 6G. However, studies based on LLMs dominate among the existing paradigms, while research on wireless BAIM (wBAIM) remains limited, and technical advances in wBAIM are still in the early stages. Furthermore, current LLM-based solutions primarily focus on a narrow set of wireless tasks, due to substantial differences in core principles and application contexts between language and wireless domains. The integration of BAIM into wireless systems has not yet reached the depth and breadth it has achieved in language and other fields. Therefore, the following sections highlight the necessity and significance of wBAIM, offering critical insights into its unique challenges and methodologies to foster a deeper integration of BAIM and wireless technologies.
@@ 3.1 Differences in Intelligence Orientation
Leveraging existing LLMs, the research value of integrating BAIM techniques and wireless networks can be initially evaluated with low resources and uncomplicated technology adaptation, which is an important motivation for for the existing researches mainly investigate wireless LLM and LLM-based wireless agent. However, there are many intuitionistic differences between language and wireless. What is the scientific nature of such differences? Can one expect to build complete universal wireless intelligence based on LLM’s transfer in the presence of such differences? This is one of the important questions to be answered in this paper.
In fact, the intelligence orientation and applications of AI technologies can be divided into two categories based on human cognition, which we term “human-centric” and “hyper-cognitive,” as illustrated in Fig. 4. The human-centric model primarily learns from human-generated data, such as images and language, and is continually refined through human feedback, such as reinforcement learning from human feedback (RLHF), ultimately aiming to replicate humanoid intelligence. In this sense, the ”teacher” of the human-centric model is humanity itself, and its ultimate goal is to replace and liberate the human being. However, due to the inherent limitations of biological intelligent agents in complex perception and intensive computation, humans lack the deep understanding and control over many natural and industrial systems. As a result, hyper-cognitive intelligence becomes essential. The hyper-cognitive model primarily learns from phenomena in natural and industrial systems, such as weather patterns and electromagnetic signals, and is optimized through system feedback, ultimately producing intelligence that surpasses the limits of human cognition. In this process, humans typically act as assistants, observing and interacting with information, while the model draws knowledge from systematic phenomena to help humans understand and manipulate complex scientific laws beyond their existing cognitive frameworks.
@@ 3.2 Scientific Significance of wBAIM
Given these substantial differences, an in-depth understanding of electromagnetic laws is essential to more freely and accurately process, represent, and utilize electromagnetic signals. This necessitates the development of a BAIM that is native to electromagnetic sciences and systems—referred to as wBAIM. However, the black-box nature of AI models, coupled with the limitations of human cognition, makes it impossible to directly assess whether a given model has truly acquired the required capabilities. As a result, implicit assessment metrics become essential. This is precisely on of the reasons why [15] defines multi-task integration, multi-scenario unification, and all-in-one scheduling as the key features of wBAIM. If a single model can be used to generalize across multiple wireless tasks and scenarios, it must inherently capture the commonalities among them—that is, the general electromagnetic laws. In other words, this is the “compression is intelligence” in the wireless context. Likely, this intelligence not only serves wireless communication systems but also provides technical supports for other electromagnetics-related fields such as remote sensing [54, 55] and radar systems [56].
@@ 4 Peculiarities of wBAIM
As discussed in Section 3, the intelligence inherent in natural language does not entirely satisfy the functional requirements of wireless systems. To fully leverage the potential of BAIM in wireless systems, it is necessary to develop wBAIM in a wireless-native manner. Before creating practical implementation methods, we must evaluate critical aspects of AI technology development, including data, models, and applications, as summarized in Fig. 5. This evaluation will help us to understand how wBAIM differs from conventional wireless models and large language models (LLMs).

### v3

# POSITION PAPER

@@ Abstract
Research on leveraging big artificial intelligence model (BAIM) technology to drive the intelligent evolution of wireless networks is emerging. However, breakthroughs in generalization brought about by BAIM techniques mainly occur in natural language processing. There is a lack of a clear technical direction on how to efficiently apply BAIM techniques to wireless systems, which typically have many additional peculiarities. To this end, this paper reviews recent research on BAIM for wireless systems and assesses the current state of the field. It then analyzes and compares the differences between language intelligence and wireless intelligence on multiple levels, including scientific foundations, core usages, and technical details. It highlights the necessity and scientific significance of developing wireless native BAIM technologies, as well as specific issues that need to be considered for technical implementation. Finally, by synthesizing the evolutionary laws of language models with the particularities of wireless systems, this paper provides several instructive methodologies for developing wireless native BAIM.
@@ 1 Introduction
Given the outstanding potential of artificial intelligence (AI) technologies in complex feature extraction, high-dimensional data representation, and adaptive decision-making, the application of AI technology to resolving problems in wireless systems has been a critical topic in wireless technology research [1, 2, 3, 4, 5, 6]. However, simple migrating or reusing successful approaches from computer vision (CV) or natural language processing (NLP) to wireless systems tends to create serious bottlenecks in terms of performance and generalizability [7]. This “incompatibility” is because AI for wireless is not directly equivalent to AI for NLP or CV. Indeed, it is necessary to incorporate electromagnetic physics and engineering methods, which makes modeling wireless AI substantially different from modeling human cognition and interaction capabilities in NLP and CV. High-quality wireless AI technologies must adapt to the characteristics of wireless data and the unique requirements of wireless systems, a concept known as becoming “wireless native.” In recent years, the incorporation of wireless native methods like physics-inspired structures [7, 8, 9, 10, 11, 12] and cross-module combination [13, 14] has significantly improved the performance and generalization of wireless AI. Meanwhile, the introduction of prior information and advanced learning architectures can reduce the demand for training data and computational resources. These fundamental advantages are of great significance for advancing the usability of wireless AI, whether in centralized or distributed usage scenarios [2].
@@
Although much progress has been made in wireless AI research, advancements have still only produced intelligence limited to specific tasks and scenarios, which creates obvious weaknesses in generalizability and stability. Meanwhile, in the NLP field, recent large language models (LLMs) have demonstrated the outstanding potential of big AI models (BAIMs) in broad generalization and high reliability. With the help of a pre-trained LLM, numerous language tasks like translation, question answering (Q&A), and summarization can be accomplished with state-of-the-art performance [15]. Moreover, the emergence of advanced LLMs like DeepSeek has significantly reduced the training and deployment costs of such large-scale AI models [16, 17, 18], further expanding the applicability of these cross-task general intelligence. Wireless AI desperately needs similar advancements in generalization. Researchers in [19] first defined wireless BAIM (wBAIM) as having the typical features of integrating multitasks, unifying multiscenarios, and all-in-one scheduling. This definition emphasizes the development of a BAIM that focuses on the physical nature of wireless data and the fundamental requirements of wireless systems. Realizing these promising visions would represent a quantum leap in functionality, universality, and reliability for wireless AI, but it requires the support of comprehensive, in-depth technical exploration. Although [19] has summarized the technical challenges and potential solutions of wBAIM, several issues almost immediately emerged with wBAIM that still require clarification and supplementation. These issues informed the research questions (RQs) for this study (see Fig. 1):
@@ 1. $\bullet$
    RQ 1: What are the most necessary and significant aspects when investigating wBAIM in the context of powerful LLMs and their transferability?
@@ 2. $\bullet$
    RQ 2: What are the technical peculiarities of wBAIM, especially in comparison to conventional wireless models and existing LLMs?
@@ 3. $\bullet$
    RQ 3: What methodologies for developing wBAIM can be derived by synthesizing the evolutionary laws of language models and the peculiarities revealed by RQ 2?
@@
To answer these RQs, this paper is organized as follows. Section 2 provides an overview of the current research landscape and recent advancements in the field. Next, in Section 3, the fundamental differences between language and wireless are analyzed, highlighting the necessity, challenges, and value of studying wBAIM. Section 4 then examines, from a system-level perspective, the unique technical considerations essential for wBAIM research. Section 5 presents several forward-looking insights into how to develop wBAIM. Finally, the conclusions are drawn in Section 6.
@@ 2 Literature Review
The success of LLMs has motivated numerous researches on BAIM for wireless to accelerate wireless systems towards ubiquitous intelligence. Existing researches can be generally categorized into the following three research paradigms: wireless LLM, LLM-based wireless agent, and wBAIM. The overall architectures of these three paradigms are shown in Fig. 2.
@@ 2.1 Wireless LLM
Wireless LLM refers to applying existing LLMs to wireless problems through reuse, fine-tuning, retrieval augmented generation (RAG) techniques, and so on. These methods translate wireless system requirements into language-based descriptions, positioning them within (or approximately within) LLMs’ generalization capabilities. Leveraging LLMs’ advanced information comprehension and vast knowledge base, initial successes have been achieved in wireless systems with protocol understanding, command interaction, and resource allocation [20, 21]. Moreover, some researchers have applied LLMs to NLP-related tasks in wireless systems, such as text categorization [22], text summarization, and Q&A [23, 24], with specific applications including the creation of AI platforms for customer interaction, virtual assistants for technicians, network diagnostic management, and telecom-standard Q&A bots [25, 26, 27]. While these studies have demonstrated the value of LLMs in wireless systems, they primarily remain within the scope of traditional NLP. Furthermore, LLMs have been applied to practical requirements in wireless systems such as power control, intrusion detection, resource allocation, and symbol demodulation [28, 29, 30, 31]. By constructing appropriate prompts, LLMs based on in-context learning can understand wireless problems to a certain extent, thus enabling initial wireless-specific functions without the need for parameter adjustments.
Meanwhile, to further enhance the applicability of LLMs in wireless communications, some researchers have begun exploring the integration of wireless domain knowledge into LLMs while focusing on the role of multimodal information [32]. These researchers have first pre-trained generative models on specific telecom knowledge datasets, followed by supervised finetuning (SFT) of the models with telecom instruction sets [33], and aligning the model’s outputs with human preferences through reinforcement learning from human feedback (RLHF) [34, 35, 36]. In addition to direct retraining, the RAG mechanism and the LangChain framework have also been identified as important supplements for a telecom knowledge base [37, 38, 39, 40, 41]. Additionally, some researchers have constructed foundational telecom text datasets to serve as material for expanding LLMs’ wireless knowledge. These include SPEC5G, a text categorization dataset derived from 3rd Generation Partnership Project (3GPP) technical reports [42]; TeleQnA, a Q&A dataset extracted from technical documents and research materials [43]; TSpec-LLM, a dataset based on 3GPP specifications that preserves table formulas [44]; Tele-Data, a LaTeX-formatted dataset obtained from sources like arXiv, 3GPP standards, and Wikipedia; and Tele-Eval, a Q&A evaluation dataset [45].
@@ 2.2 LLM-Based Wireless Agent
Although LLM-based methods have been somewhat explored, the current technical progress primarily focuses on instruction design, resource scheduling, and a few simple wireless problems. These approaches still lack effective solutions for numerous tough wireless challenges, such as scenario sensing and high-rate transmission. This limitation arises from two primary factors: firstly, linguistic expressions are frequently insufficient in precisely characterizing multimodal signals and their intricate correlations; secondly, the language intelligence derived from textual training lacks adequate knowledge reserves to comprehend the data structures and underlying mechanisms inherent in wireless data. Therefore, instead of embracing LLMs, some researchers have explored BAIM in a wireless native way. Wireless native AI refers to an intelligent framework based on regular patterns of AI technologies, such as deep neural networks, while incorporating the peculiarities of wireless systems. It tailors model design to the nature of wireless data and system requirements, and adapts through re-learning on wireless data. Conventional wireless AI has also followed this approach in many studies. For example, in model design, researchers initially leveraged the numerical characteristics of wireless data, such as exploiting channel sparsity in the angle-delay domain and transforming it into the angle domain to make it more image-like [51, 52]. Later, researchers delved deeper into the physical mechanisms underlying the numerical characteristics and employed more fundamental, physics-inspired techniques to guide model optimization [7, 8, 9, 10, 12]. In terms of architecture, some cross-module [13, 14] and cross-layer [53] joint techniques have also been developed to facilitate applications in wireless systems.
However, given that wBAIM has a broader functional scope, relying solely on existing wireless AI technologies is insufficient. Therefore, in [19], the authors first outlined three key generalizability indicators of wBAIM: multi-task integration, multi-scenario unification, and all-in-one scheduling. Meanwhile, [19] also provides a demonstrative case study where a CMixer [11] model is trained across multiple scenarios to simultaneously accomplish high-dimensional channel estimation and channel compression feedback tasks, illustrating the feasibility and significance of wBAIM. Building on this basis, [54] introduced SWTCAN, a network based on the advanced swin transformer, which integrates downlink pilot patterns, channel estimation, and uplink channel feedback into a joint design. [55] proposed Csi-LLM, a downlink channel prediction model based on GPT2, which supports variable-length inputs and enables continuous prediction. Similarly, in [56], a pre-training and fine-tuning based channel prediction method was presented, enhancing performance in few-shot learning. Meanwhile, [57] evaluated the similarity and diversity across multi-scenario datasets, offering valuable insights into developing multi-scenario generalization techniques from a data-based perspective.
@@ 2.4 Summary of the Literature
Fig. 3 presents statistical information on the 36 publications related to BAIM in wireless systems included in this paper. It shows that research interest in BAIM for wireless is growing, and there is a growing consensus on the importance of BAIM techniques for the evolution of 6G. However, studies based on LLMs dominate among the existing paradigms, while research on wBAIM remains limited, and technical advances in wBAIM are still in the early stages. Furthermore, current LLM-based solutions primarily focus on a narrow set of wireless tasks, due to substantial differences in core principles and application contexts between language and wireless domains. The integration of BAIM into wireless systems has not yet reached the depth and breadth it has achieved in language and other fields. Therefore, the following sections highlight the necessity and significance of wBAIM, offering critical insights into its unique challenges and methodologies to foster a deeper integration of BAIM and wireless technologies.
@@ 3.1 Differences in Intelligence Orientation
Leveraging existing LLMs, the research value of integrating BAIM techniques and wireless networks can be initially evaluated with low resources and uncomplicated technology adaptation, which is an important motivation for why existing research primarily investigates wireless LLM and LLM-based wireless agent. However, there are many intuitionistic differences between language and wireless. What is the scientific nature of such differences? Can one expect to build complete universal wireless intelligence based on LLM’s transfer in the presence of such differences? This is one of the important questions to be answered in this paper.
In fact, the intelligence orientation and applications of AI technologies can be divided into two categories based on human cognition, which we term “human-centric” and “hyper-cognitive,” as illustrated in Fig. 4. The human-centric model primarily learns from human-generated data, such as images and language, and is continually refined through human feedback, such as reinforcement learning from human feedback (RLHF), ultimately aiming to replicate humanoid intelligence. In this sense, the ”teacher” of the human-centric model is humanity itself, and its ultimate goal is to replace and liberate humans. However, due to the inherent limitations of biological intelligent agents in complex perception and intensive computation, humans lack the deep understanding and control over many natural and industrial systems. As a result, hyper-cognitive intelligence becomes essential. The hyper-cognitive model primarily learns from phenomena in natural and industrial systems, such as weather patterns and electromagnetic signals, and is optimized through system feedback, ultimately producing intelligence that surpasses the limits of human cognition. In this process, humans typically act as assistants, observing and interacting with information, while the model draws knowledge from systematic phenomena to help humans understand and manipulate complex scientific laws beyond their existing cognitive frameworks.
@@ 3.2 Scientific Significance of wBAIM
Given these substantial differences, an in-depth understanding of electromagnetic laws is essential to more freely and accurately process, represent, and utilize electromagnetic signals. This necessitates the development of a BAIM that is native to electromagnetic sciences and systems—referred to as wBAIM. However, the black-box nature of AI models, coupled with the limitations of human cognition, makes it impossible to directly assess whether a given model has truly acquired the required capabilities. As a result, implicit assessment metrics become essential. This is precisely one of the reasons why [19] defines multi-task integration, multi-scenario unification, and all-in-one scheduling as the key features of wBAIM. If a single model can be used to generalize across multiple wireless tasks and scenarios, it must inherently capture the commonalities among them—that is, the general electromagnetic laws. In other words, this is the “compression is intelligence” in the wireless context. Likely, this intelligence not only serves wireless communication systems but also provides technical supports for other electromagnetics-related fields such as remote sensing [58, 59] and radar systems [60].
@@ 4 Peculiarities of wBAIM
As discussed in Section 3, the intelligence inherent in natural language does not entirely satisfy the functional requirements of wireless systems. To fully leverage the potential of BAIM in wireless systems, it is necessary to develop wBAIM in a wireless native manner. Before creating practical implementation methods, we must evaluate critical aspects of AI technology development, including data, models, and applications, as summarized in Fig. 5. This evaluation will help us to understand how wBAIM differs from conventional wireless models and LLMs.